What frequency of sound should a killer whale emit to ensure that it receives an echo from a seal that is 1.5 m in length?
Answer #1

Solution) L = 1.5 m

f = ?

V = (D/T)

But (1/T) = f

V = Df

V = 1531 m/s in sea water

D = 2L = 2×1.5 = 3 m

f = (V)/(D)

f = (1531)/(3)

f = 510.3 Hz